OpenStudy (anonymous):

Jaclyn is selling cookies for her club. She sells 2 boxes for $5. What unit rate expresses the cost per box? $1 per box $1.50 per box $2 per box $2.50 per box

OpenStudy (anonymous):

i dont get this omo

OpenStudy (anonymous):

what does each box of cookies cost is what the question is asking

OpenStudy (anonymous):

if you buy 2 chocolates for $10, what is the cost of 1 chocolate??

OpenStudy (anonymous):

@harikat 5$

OpenStudy (anonymous):

Spelled ur name wrong sry

OpenStudy (anonymous):

how did you get this answer??

OpenStudy (anonymous):

what would one of the boxes of cookies cost?

OpenStudy (anonymous):

10/2=5

OpenStudy (anonymous):

Ok 2 boxes$5 and 1 box$?

OpenStudy (anonymous):

correct... you have to do the same to find the cost of one box

OpenStudy (anonymous):

I don't no this

OpenStudy (anonymous):

it is D 2.50 per box

OpenStudy (anonymous):

it is the unit rate

OpenStudy (anonymous):

yeah i divided 5/2 and i got that y did they say 10 for 2

OpenStudy (anonymous):

uno is 1 is spanish unit is 1

OpenStudy (anonymous):

Idk they not smart

OpenStudy (anonymous):

can I have medal?

OpenStudy (anonymous):

they gonna copy me now

OpenStudy (anonymous):

and thx

OpenStudy (anonymous):

just to make it easier for you to understand...

OpenStudy (anonymous):

divide it by the number

OpenStudy (anonymous):

hark im in middle school :3 I get what you were doin but i could understand if you put it in diff terms

OpenStudy (anonymous):

you can do $2 1/2 I guess

OpenStudy (anonymous):

and I am in middle school too

OpenStudy (anonymous):

here the requested rate r is given by the subsequent computation: r=5/2= 2.5 I think :D

OpenStudy (anonymous):

studygirl trust me you copied me when you first were here saying idk the answer to this stuff

OpenStudy (anonymous):

try experimenting with this

OpenStudy (anonymous):

trying to hit 50 today